OpenStudy (anonymous):

What's the inverse of a 2x4 matrix?

OpenStudy (unklerhaukus):

can a matrix that is not square have an inverse ?

OpenStudy (unklerhaukus):

remember \[\textbf A \textbf A^{-1}=\textbf I=\textbf A^{-1} \textbf A\]

OpenStudy (zarkon):

it could have a right invers

OpenStudy (zarkon):

*inverse

OpenStudy (anonymous):

How then is it possible to use matrices more generally in simultaneous equations (as I gather they are) without being limited to overly specific situations?

OpenStudy (unklerhaukus):

what is a right inverse @Zarkon ?

OpenStudy (amistre64):

a matrix that is multiplied in the right sid ethat make an indentity matrix

OpenStudy (zarkon):

for an \(m\times n\) matrix A (\(m<n\)) of full rank you can use \[A^T \left(A A^T\right)^{-1}\]

OpenStudy (anonymous):

Non-square matrices (m-by-n matrices for which m ≠ n) do not have an inverse. However, in some cases such a matrix may have a left inverse or right inverse. If A is m-by-n and the rank of A is equal to n, then A has a left inverse: an n-by-m matrix B such that BA = I. If A has rank m, then it has a right inverse: an n-by-m matrix B such that AB = I.

OpenStudy (zarkon):

then \[AA^T \left(A A^T\right)^{-1}=I\]

OpenStudy (unklerhaukus):

so there are two solutions to this question,